# spacing=nonfrench

### v1

@@ V-A Experimental Settings
Dataset: We use the Waymo Open Motion Dataset (WOMD) [53] for which 1.1 seconds is used to predict the 8-second future trajectory of target agents. For the robustness experiments, similar to [5], we modify the set of target agents to have: (i) PlusAV setting for which the autonomous vehicle (AV) is added to the set of target agents, and (ii) OnlyAV setting, where the AV is the only target agent.
@@
Evaluation Metrics. We use the standard benchmark metrics of WOMD and INTERACTION including minADE, minFDE, mAP, and Miss Rate (MR). Following the evaluation protocol in [5], we also assess the model’s robustness to causality-based perturbation by reporting $\Delta$-metrics given by $\frac{1}{N}\sum_{i=1}^{N}\left|m^{i}_{\textup{Original}}-m^{i}_{\textup{Perturbed}}\right|$.
@@ V-B Quantitative Analysis
Robustness Experiments. To study the robustness and generalizability of the CRiTIC model under causality-based domain shifts, we use scene perturbation tests used in [5]. In Tab. I, we present the result for RemoveNonCausal perturbation, where all agents tagged as non-causal are removed. Evaluating based on the relative performance drop rate metric, $\frac{\Delta\textup{minADE}}{\textup{minADE}_{\textup{Original}}}$, our model shows significant improvement by up to 54% and 53% compared to the next best model for both the OnlyAV and PlusAV settings, respectively. As expected, the CRiTIC model has less performance drop for RemoveNonCausal perturbations as its attention is directed towards causal agents, and it is less reliant on non-causal agents.

### v2

@@ V-A Experimental Settings
Dataset: We use the Waymo Open Motion Dataset (WOMD) [53] for which 1.1 seconds is used to predict the 8-second future trajectory of target agents. For the robustness experiments, similar to [5], we modify the set of target agents to have: (i) PlusAV setting for which the autonomous vehicle (AV) (which is specified in the dataset) is added to the set of target agents, and (ii) OnlyAV setting, where the AV is the only target agent.
@@
Evaluation Metrics. We use the standard benchmark metrics of WOMD and INTERACTION including minADE, minFDE, mAP, and Miss Rate (MR) [53]. Following the evaluation protocol in [5], we also assess the model’s robustness to causality-based perturbation by reporting $\Delta$-metrics given by $\frac{1}{N}\sum_{i=1}^{N}\left|m^{i}_{\textup{Original}}-m^{i}_{\textup{Perturbed}}\right|$.
@@ V-B Quantitative Analysis
Robustness Experiments. To study the robustness and generalization of the CRiTIC model under causality-based domain shifts, we use scene perturbation tests with the causal agent annotation by Sun et al. [5]. In Tab. I, we present the result for RemoveNonCausal perturbation, where all agents tagged as non-causal are removed. Evaluating based on the relative performance drop rate metric, $\frac{\Delta\textup{minADE}}{\textup{minADE}_{\textup{Original}}}$, our model shows significant improvement by up to 54% and 53% compared to the next best model for both the OnlyAV and PlusAV settings, respectively. As expected, the CRiTIC model has less performance drop for RemoveNonCausal perturbations as its attention is directed towards causal agents, and it is less reliant on non-causal agents. Note that, we use the minADE metric in the robustness experiments following [5], however, we should note that it is not the primary ranking metric in Waymo’s prediction challenge.